OpenStudy (anonymous):

find the slope of the tangent line to the curve -1x^2-2xy-3xy^3=-185 at point (-1,4)

OpenStudy (anonymous):

sorry i mean at point (-2,-4)

OpenStudy (anonymous):

The derivative if this function is-2x-2y-2xy'-3y^3-3x3y^2y'

OpenStudy (anonymous):

Y'=...... Make the (-2,-4)in

OpenStudy (anonymous):

oh crap i wrote the equation wrong... sorry it's actually -1x^2-4xy-2y^3=92

OpenStudy (anonymous):

The derivative if this function is-2x-2y-2xy'-3y^3-3x3y^2y'=-185.so y'=. And then make (-2,-4)in.

OpenStudy (anonymous):

ok so if we're taking the derivative of this equation -1x^2-4xy-2y^3=92

OpenStudy (anonymous):

would it be -2x-4xy'=4y-6y^2y'=0

OpenStudy (anonymous):

No, it is a implicit function.

OpenStudy (anonymous):

yes it is, i took the derivative and that's what i got, idk if it's right tho

OpenStudy (anonymous):

So we can make it-------- -2x-4y-4xy'-6y^2y'=92. X=-2. Y=-4 So. 4+16+8y'-96y'=92 so y'=72/-88=-9/11

OpenStudy (anonymous):

Maybe I am wrong. It is too hard to write on iPad

OpenStudy (anonymous):

Sorry i am wrong. I forget to make 92 into 0

OpenStudy (anonymous):

yea so it would be -2x-4xy'=4y-6y^2=0

OpenStudy (anonymous):

sorry typo... i meant -2x-4xy'+4y-6y^2y'=0